# Conditionally valid Probabilistic Conformal Prediction

### v1

@@ 5 Conclusion
We address the challenge of conditional coverage in conformal predictions, and overcome previous negative results by assuming the knowledge of a good estimator of $\textup{P}_{Y\mid X}$. We develop a general framework for constructing prediction sets, which encompasses many existing conformal methods and refine their confidence sets using a probabilistic estimator of $Y\mid X$. Our key contributions include theoretical studies of the conditional validity, holding significant implications for the development of future methods.
Acknowledgments and Disclosure of Funding
@@ Appendix A Additional results and calculations
In this section, we analyze the theoretical results of Section 3. First, let’s recall the definition of the quantile function for any distribution $\mu$ living in $\mathbb{R}$. For any $\alpha\in(0,1)$, the quantile $Q_{1-\alpha}(\mu)$ is defined by
$Q_{1-\alpha}(\mu)=\inf\left\{{t\in\mathbb{R}\colon\mu((-\infty,t])\geq 1-\alpha}\right\}.$
@@
$\tau_{x,z}=\inf\left\{{\tau\in\mathbb{R}\colon\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau}(\varphi)))\geq 1-\alpha}\right\},$ (7)
$\lambda_{x,y,z}=\inf\left\{{\lambda\in\mathbb{R}\colon y\in\mathcal{R}_{z}(x;\lambda)}\right\},$
@@
$\mathcal{C}_{\alpha}(x)=\mathcal{R}_{z}\left({x;f_{\tau_{x,z}}\bigl{(}Q_{1-\alpha}(\mu)\bigr{)}}\right),$
where the distribution $\mu$ is defined as
$\mu=\frac{1}{n+1}\sum_{k=1}^{n}\delta_{f_{\bar{\tau}_{k}}^{-1}(\bar{\lambda}_{k})}+\frac{1}{n+1}\delta_{\infty}.$
@@
Assume 1 hold. For any $(x,y,z)\in\mathbb{R}^{d}\times\mathcal{Y}\times\mathcal{Z}$, $\lambda_{x,y,z}$ exists in $\mathbb{R}$, and we have $y\in\mathcal{R}_{z}(x;\lambda_{x,y,z})$.
@@
Let $(x,y,z)\in\mathbb{R}^{d}\times\mathcal{Y}\times\mathcal{Z}$ be fixed. Since $\cap_{t\in\mathbb{R}}\mathcal{R}_{z}(x;t)=\emptyset$ and $\cup_{t\in\mathbb{R}}\mathcal{R}_{z}(x;t)=\mathcal{Y}$, we deduce the existence of $t_{0}$ and $t_{1}$ such that $y\notin\mathcal{R}_{z}(x;t_{0})$ and $y\in\mathcal{R}_{z}(x;t_{1})$. Therefore, $\{t\in\mathbb{R}\colon y\in\mathcal{R}_{z}(x;t)\}$ is non-empty and lower-bounded by $t_{0}$. Thus, the infimum $\lambda_{x,y,z}$ exists. Now, let’s prove that $y\in\mathcal{R}_{z}(x;\lambda_{x,y,z})$. Since $\lambda_{x,y,z}=\inf\{t\in\mathbb{R}\colon y\in\mathcal{R}_{z}(x;t)\}$, we deduce the existence of a decreasing sequence $\{\lambda_{n}\}_{n\in\mathbb{N}}$ such that $y\in\mathcal{R}_{z}(x;\lambda_{n})$ and $\lim_{n\to\infty}\lambda_{n}=\lambda_{x,y,z}$. By definition of $\{\lambda_{n}\}_{n\in\mathbb{N}}$, we have $y\in\cap_{n\in\mathbb{N}}\mathcal{R}_{z}(x;\lambda_{n})$. However, using 1, remark that
@@
$1-\alpha\leq\mathbb{P}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})}\right)<1-\alpha+\frac{1}{n+1}.$ (8)
@@
By definition, we have
$\mathbb{P}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})}\right)$ $=\mathbb{P}\left({Y_{n+1}\in\mathcal{R}_{Z_{n+1}}\left({X_{n+1},f_{\bar{\tau}_{n+1}}(Q_{1-\alpha}(\mu))}\right)}\right)$
$=\mathbb{P}\left({\lambda_{n+1}\leq f_{\bar{\tau}_{n+1}}(Q_{1-\alpha}(\mu))}\right).$
@@
$\mathbb{P}\left({\lambda_{n+1}\leq f_{\bar{\tau}_{n+1}}(Q_{1-\alpha}(\mu))}\right)=\mathbb{P}\left({f_{\bar{\tau}_{n+1}}^{-1}(\lambda_{n+1})\leq Q_{1-\alpha}(\mu)}\right).$
@@
$\mathbb{P}\left({V_{n+1}\leq Q_{1-\alpha}\left({\sum_{k=1}^{n}\frac{\delta_{V_{k}}}{n+1}+\frac{\delta_{\infty}}{n+1}}\right)}\right)=\mathbb{P}\left({V_{n+1}\leq Q_{1-\alpha}\left({\sum_{k=1}^{n+1}\frac{\delta_{V_{k}}}{n+1}}\right)}\right)\\ =\frac{1}{n+1}\sum_{k=1}^{n+1}\mathbb{E}\left[{\mathds{1}_{V_{k}}\leq Q_{1-\alpha}\left({\frac{1}{n+1}\sum_{k=1}^{n+1}\delta_{V_{k}}}\right)}\right]\\ =\mathbb{E}\left[{\mathbb{E}\left[{\mathds{1}_{V_{I}}\leq Q_{1-\alpha}\left({\frac{1}{n+1}\sum_{k=1}^{n+1}\delta_{V_{k}}}\right)\,\bigg{|}\,V_{1},\ldots,V_{n+1}}\right]}\right],$
@@
$\mathbb{P}\left({f_{\bar{\tau}_{n+1}}^{-1}(\lambda_{n+1})\leq Q_{1-\alpha}(\mu)}\right)<1-\alpha+\frac{1}{n+1}.$
@@
Assume 1-2 hold, and let $\alpha\in(0,1$), $x\in\mathbb{R}^{d}$, $z\in\mathcal{Z}$. If $\Pi_{Y\mid X=x}$ is a probability measure, then $\tau_{x,z}$ is defined in $\mathbb{R}$ and $\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau_{x,z}}(\varphi)))\geq 1-\alpha$.
@@
$\lim\limits_{\tau\to+\infty}\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau}(\varphi)))$ $=\Pi_{Y\mid X=x}\left({\cup_{\tau\in\mathbb{R}}\mathcal{R}_{z}(x;f_{\tau}(\varphi))}\right)$
$=\Pi_{Y\mid X=x}\left({\cup_{t\in\mathbb{R}}\mathcal{R}_{z}(x;t)}\right)=1.$
The previous equality shows the existence of $\tau\in\mathbb{R}$ such that $\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau}(\varphi)))\geq 1-\alpha$. Therefore $\{\tau\in\mathbb{R}\colon\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau}(\varphi)))\geq 1-\alpha\}$ is non-empty. This proves the existence of $\tau_{x,z}=\inf\{\tau\in\mathbb{R}\colon\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau}(\varphi)))\geq 1-\alpha\}$ in ${}\mathbb{R}\cup\{-\infty\}$. However, $\tau_{x,z}>-\infty$, otherwise we would have
$01-\alpha\leq\lim\limits_{\tau\to-\infty}\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau}(\varphi)))=\Pi_{Y\mid X=x}\left({\cap_{t\in\mathbb{R}}\mathcal{R}_{z}(x;t)}\right)=\Pi_{Y\mid X=x}(\emptyset)=$ 0 .
Therefore, we deduce that $\tau_{x,z}\in\mathbb{R}$. Lastly, remark that
@@
Now, we prove Theorem 3.2. This result guarantees that the conditional confidence intervals constructed by our method approximately satisfy the desired coverage of $1-\alpha$.
@@
$\mathbb{P}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})\mid X_{n+1}=x,Z_{n+1}=z}\right)\geq 1-\alpha-\mathrm{d}_{\mathrm{TV}}(\textup{P}_{Y\mid X=x};\Pi_{Y\mid X=x})\\ -\mathbb{P}\left({Q_{1-\alpha}(\mu)<f_{\tau_{x,z}}^{-1}(\lambda_{x,Y_{n+1},z})\leq\varphi\,|\,X_{n+1}=x,Z_{n+1}=z}\right).$
@@
First, recall that $\mathcal{C}_{\alpha}(x)$ is given in (4), and $\lambda_{x,Y_{n+1},z}$ is defined in (7). Applying Lemma A.1, we know that $\lambda_{x,Y_{n+1},z}$ is defined in $\mathbb{R}$, and also that $Y_{n+1}\in\mathcal{R}_{z}(x;\lambda_{x,Y_{n+1},z})$. Hence, it holds
$\mathbb{P}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})\,|\,X_{n+1}=x,Z_{n+1}=z}\right)$
$=\mathbb{P}\left({Y_{n+1}\in\mathcal{R}_{z}\left({x;f_{\tau_{x,z}}(Q_{1-\alpha}(\mu))}\right)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)$
$=\mathbb{P}\left({\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}\left({Q_{1-\alpha}(\mu)}\right)\,|\,X_{n+1}=x,Z_{n+1}=z}\right).$
Let’s introduce the term $\mathbb{P}(\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}(\varphi)\,|\,X_{n+1}=x,Z_{n+1}=z)$ as follows:
$\mathbb{P}\left({\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}\left({Q_{1-\alpha}(\mu)}\right)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ =\mathbb{P}\left({\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}\left({Q_{1-\alpha}(\mu)}\right)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ \pm\mathbb{P}\left({\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}(\varphi)\,|\,X_{n+1}=x,Z_{n+1}=z}\right).$ (9)
@@
$A=\mathbb{P}\left({f_{\tau_{x,z}}^{-1}(\lambda_{x,Y_{n+1},z})\leq Q_{1-\alpha}(\mu)<\varphi\,|\,X_{n+1}=x,Z_{n+1}=z}\right),$
$B=\mathbb{P}\left({f_{\tau_{x,z}}^{-1}(\lambda_{x,Y_{n+1},z})\leq\varphi\leq Q_{1-\alpha}(\mu)\,|\,X_{n+1}=x,Z_{n+1}=z}\right).$
@@
$\mathbb{P}\left({\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}\left({Q_{1-\alpha}(\mu)}\right)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ =A+B+\mathbb{P}\left({\varphi<f_{\tau_{x,z}}^{-1}\left({\lambda_{x,Y_{n+1},z}}\right)\leq Q_{1-\alpha}(\mu)\,|\,X_{n+1}=x,Z_{n+1}=z}\right),$
@@
$\mathbb{P}\left({\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}(\varphi)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ =A+B+\mathbb{P}\left({Q_{1-\alpha}(\mu)<f_{\tau_{x,z}}^{-1}\left({\lambda_{x,Y_{n+1},z}}\right)\leq\varphi\,|\,X_{n+1}=x,Z_{n+1}=z}\right).$
@@
$\mathbb{P}\left({\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}\left({Q_{1-\alpha}(\mu)}\right)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ -\mathbb{P}\left({\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}(\varphi)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ =\mathbb{P}\left({\varphi<f_{\tau_{x,z}}^{-1}\left({\lambda_{x,Y_{n+1},z}}\right)\leq Q_{1-\alpha}(\mu)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ -\mathbb{P}\left({Q_{1-\alpha}(\mu)<f_{\tau_{x,z}}^{-1}\left({\lambda_{x,Y_{n+1},z}}\right)\leq\varphi\,|\,X_{n+1}=x,Z_{n+1}=z}\right).$ (10)
@@
$\mathbb{P}\left({\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}(\varphi)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ \geq\mathbb{P}\left({\lambda_{x,\hat{Y}_{n+1},z}\leq f_{\tau_{x,z}}(\varphi)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)-\mathrm{d}_{\mathrm{TV}}(\textup{P}_{Y\mid X=x};\Pi_{Y\mid X=x}).$
@@
$\mathbb{P}\left({\lambda_{x,\hat{Y}_{n+1},z}\leq f_{\tau_{x,z}}(\varphi)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ =\mathbb{P}\left({\hat{Y}_{n+1}\in\left\{{y\in\mathcal{Y}\colon\lambda_{x,y,z}\leq f_{\tau_{x,z}}(\varphi)}\right\}\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ =\mathbb{P}\left({\hat{Y}_{n+1}\in\mathcal{R}_{z}(x;f_{\tau_{x,z}}(\varphi))\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ =\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau_{x,z}}(\varphi)))\geq 1-\alpha.$

### v2

@@ 5 Conclusion
We address the challenge of conditional coverage in CP, and overcome previous negative results by assuming the knowledge of a good estimator of $\textup{P}_{Y\mid X}$. Our proposed mechanism conformalized the conditional estimator $\Pi_{Y\mid X}$ to ensure marginal validity while maintaining similar conditional coverage guarantees. Specifically, if experts can provide an accurate conditional estimator, our algorithm CP2 generates nearly conditionally valid multidimensional prediction sets. This approach offers a practical solution for tackling heteroscedasticity in various machine learning applications.
Acknowledgments
@@ Appendix A Additional Results and Calculations
In this section, we analyze the theoretical results of Section 3. First, let’s recall the definition of the quantile function for any distribution $\mu_{n}$ living in $\mathbb{R}$. For any $\alpha\in(0,1)$, the quantile $Q_{1-\alpha}(\mu_{n})$ is defined by
$Q_{1-\alpha}(\mu_{n})=\inf\left\{{t\in\mathbb{R}\colon\mu_{n}((-\infty,t])\geq 1-\alpha}\right\}.$
@@
$\tau_{x,z}=\inf\left\{{\tau\in\mathsf{T}\colon\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau}(\varphi)))\geq 1-\alpha}\right\},$ (7)
$\lambda_{x,y,z}=\inf\left\{{\lambda\in\mathsf{T}\colon y\in\mathcal{R}_{z}(x;\lambda)}\right\},$
@@
$\mathcal{C}_{\alpha}(x)=\mathcal{R}_{z}\left({x;f_{\tau_{x,z}}\bigl{(}Q_{1-\alpha}(\mu_{n})\bigr{)}}\right),$
where the distribution $\mu_{n}$ is defined as
$\mu_{n}=\frac{1}{n+1}\sum_{k=1}^{n}\delta_{f_{\bar{\tau}_{k}}^{-1}(\bar{\lambda}_{k})}+\frac{1}{n+1}\delta_{\infty}.$
@@
Assume 1 hold. For any $(x,y,z)\in\mathbb{R}^{d}\times\mathcal{Y}\times\mathcal{Z}$, $\lambda_{x,y,z}$ exists in $\mathsf{T}$, and we have $y\in\mathcal{R}_{z}(x;\lambda_{x,y,z})$.
@@
Let $(x,y,z)\in\mathbb{R}^{d}\times\mathcal{Y}\times\mathcal{Z}$ be fixed. Since $\cap_{t\in\mathsf{T}}\mathcal{R}_{z}(x;t)=\emptyset$ and $\cup_{t\in\mathsf{T}}\mathcal{R}_{z}(x;t)=\mathcal{Y}$, we deduce the existence of $t_{0}$ and $t_{1}$ such that $y\notin\mathcal{R}_{z}(x;t_{0})$ and $y\in\mathcal{R}_{z}(x;t_{1})$. Therefore, $\{t\in\mathsf{T}\colon y\in\mathcal{R}_{z}(x;t)\}$ is non-empty and lower-bounded by $t_{0}$. Thus, the infimum $\lambda_{x,y,z}$ exists. Now, let’s prove that $y\in\mathcal{R}_{z}(x;\lambda_{x,y,z})$. Since $\lambda_{x,y,z}=\inf\{t\in\mathsf{T}\colon y\in\mathcal{R}_{z}(x;t)\}$, we deduce the existence of a decreasing sequence $\{\lambda_{n}\}_{n\in\mathbb{N}}$ such that $y\in\mathcal{R}_{z}(x;\lambda_{n})$ and $\lim_{n\to\infty}\lambda_{n}=\lambda_{x,y,z}$. By definition of $\{\lambda_{n}\}_{n\in\mathbb{N}}$, we have $y\in\cap_{n\in\mathbb{N}}\mathcal{R}_{z}(x;\lambda_{n})$. However, using 1, remark that
@@
$1-\alpha\leq\mathbb{P}^{\mathcal{T}}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})}\right)<1-\alpha+\frac{1}{n+1}.$ (8)
@@
Using Lemma A.1, we have
$\mathbb{P}^{\mathcal{T}}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})}\right)$ $=\mathbb{P}^{\mathcal{T}}\left({Y_{n+1}\in\mathcal{R}_{Z_{n+1}}\left({X_{n+1},f_{\bar{\tau}_{n+1}}(Q_{1-\alpha}(\mu_{n}))}\right)}\right)$
$=\mathbb{P}^{\mathcal{T}}\left({\lambda_{n+1}\leq f_{\bar{\tau}_{n+1}}(Q_{1-\alpha}(\mu_{n}))}\right).$
@@
$\mathbb{P}^{\mathcal{T}}\left({\lambda_{n+1}\leq f_{\bar{\tau}_{n+1}}(Q_{1-\alpha}(\mu_{n}))}\right)=\mathbb{P}^{\mathcal{T}}\left({f_{\bar{\tau}_{n+1}}^{-1}(\lambda_{n+1})\leq Q_{1-\alpha}(\mu_{n})}\right).$
@@
$\mathbb{P}^{\mathcal{T}}\left({V_{n+1}\leq Q_{1-\alpha}\left({\sum_{k=1}^{n}\frac{\delta_{V_{k}}}{n+1}+\frac{\delta_{\infty}}{n+1}}\right)}\right)=\mathbb{P}^{\mathcal{T}}\left({V_{n+1}\leq Q_{1-\alpha}\left({\sum_{k=1}^{n+1}\frac{\delta_{V_{k}}}{n+1}}\right)}\right)\\ =\frac{1}{n+1}\sum_{k=1}^{n+1}\mathbb{E}^{\mathcal{T}}\left[{\mathds{1}_{V_{k}}\leq Q_{1-\alpha}\left({\frac{1}{n+1}\sum_{k=1}^{n+1}\delta_{V_{k}}}\right)}\right]\\ =\mathbb{E}^{\mathcal{T}}\left[{\mathbb{E}^{\mathcal{T}}\left[{\mathds{1}_{V_{I}}\leq Q_{1-\alpha}\left({\frac{1}{n+1}\sum_{k=1}^{n+1}\delta_{V_{k}}}\right)\,\bigg{|}\,V_{1},\ldots,V_{n+1}}\right]}\right],$
@@
$\mathbb{P}^{\mathcal{T}}\left({f_{\bar{\tau}_{n+1}}^{-1}(\lambda_{n+1})\leq Q_{1-\alpha}(\mu_{n})}\right)<1-\alpha+\frac{1}{n+1}.$
@@
Assume 1-2 hold, and let $\alpha\in(0,1$), $x\in\mathbb{R}^{d}$, $z\in\mathcal{Z}$. If $\Pi_{Y\mid X=x}$ is a probability measure, then $\tau_{x,z}$ is defined in $\mathsf{T}$ and $\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau_{x,z}}(\varphi)))\geq 1-\alpha$.
@@
$\sup_{\tau\in\mathsf{T}}\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau}(\varphi)))$ $=\Pi_{Y\mid X=x}\left({\cup_{\tau\in\mathsf{T}}\mathcal{R}_{z}(x;f_{\tau}(\varphi))}\right)$
$=\Pi_{Y\mid X=x}\left({\cup_{t\in\mathsf{T}}\mathcal{R}_{z}(x;t)}\right)=1.$
The previous equality shows the existence of $\tau\in\mathsf{T}$ such that $\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau}(\varphi)))\geq 1-\alpha$. Therefore $\{\tau\in\mathsf{T}\colon\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau}(\varphi)))\geq 1-\alpha\}$ is non-empty. This proves the existence of $\tau_{x,z}=\inf\{\tau\in\mathsf{T}\colon\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau}(\varphi)))\geq 1-\alpha\}$ in $\mathsf{T}\cup\{-\infty\}$. Moreover, $\tau_{x,z}>-\infty$, otherwise we would have
$01-\alpha\leq\inf_{\tau\in\mathsf{T}}\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau}(\varphi)))=\Pi_{Y\mid X=x}\left({\cap_{t\in\mathsf{T}}\mathcal{R}_{z}(x;t)}\right)=$ 0 .
Therefore, we deduce that $\tau_{x,z}\in\mathsf{T}$. Lastly, remark that
@@
Now, we prove Theorem 3.2. This result guarantees that the conditional confidence intervals constructed by our method approximately satisfy the desired coverage of $1-\alpha$. Given $(x,y)\in\mathbb{R}^{d}\times\mathcal{Z}$, let’s introduce
$p_{n+1}^{(x,z)}=\mathbb{P}^{\mathcal{T}}\left({Q_{1-\alpha}(\mu_{n})<f_{\tau_{x,z}}^{-1}\left({\lambda_{x,Y_{n+1},z}}\right)\leq\varphi\,|\,X_{n+1}=x,Z_{n+1}=z}\right),$
$q_{n+1}^{(x,z)}=\mathbb{P}^{\mathcal{T}}\left({\varphi<f_{\tau_{x,z}}^{-1}\left({\lambda_{x,Y_{n+1},z}}\right)\leq Q_{1-\alpha}(\mu_{n})\,|\,X_{n+1}=x,Z_{n+1}=z}\right).$
@@
$1-\alpha-\mathrm{d}_{\mathrm{TV}}(\textup{P}_{Y\mid X=x};\Pi_{Y\mid X=x})-p_{n+1}^{(x,z)}\leq\mathbb{P}^{\mathcal{T}}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})\mid X_{n+1}=x,Z_{n+1}=z}\right)\\ \leq\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau_{x,z}}(\varphi)))+\mathrm{d}_{\mathrm{TV}}(\textup{P}_{Y\mid X=x};\Pi_{Y\mid X=x})+q_{n+1}^{(x,z)}.$
@@
First, recall that $\mathcal{C}_{\alpha}(x)$ is given in (4), and $\lambda_{x,Y_{n+1},z}$ is defined in (7). Applying Lemma A.1, we know that $\lambda_{x,Y_{n+1},z}$ is defined in $\mathsf{T}$, and also that $Y_{n+1}\in\mathcal{R}_{z}(x;\lambda_{x,Y_{n+1},z})$. Hence, it holds
$\mathbb{P}^{\mathcal{T}}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ =\mathbb{P}^{\mathcal{T}}\left({Y_{n+1}\in\mathcal{R}_{z}\left({x;f_{\tau_{x,z}}(Q_{1-\alpha}(\mu_{n}))}\right)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ =\mathbb{P}^{\mathcal{T}}\left({\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}\left({Q_{1-\alpha}(\mu_{n})}\right)\,|\,X_{n+1}=x,Z_{n+1}=z}\right).$ (9)
Let’s introduce the term $\mathbb{P}^{\mathcal{T}}(\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}(\varphi)\,|\,X_{n+1}=x,Z_{n+1}=z)$ as follows:
$\mathbb{P}^{\mathcal{T}}\left({\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}\left({Q_{1-\alpha}(\mu_{n})}\right)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ =\mathbb{P}^{\mathcal{T}}\left({\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}\left({Q_{1-\alpha}(\mu_{n})}\right)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ \pm\mathbb{P}^{\mathcal{T}}\left({\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}(\varphi)\,|\,X_{n+1}=x,Z_{n+1}=z}\right).$ (10)
@@
$A=\mathbb{P}^{\mathcal{T}}\left({f_{\tau_{x,z}}^{-1}(\lambda_{x,Y_{n+1},z})\leq Q_{1-\alpha}(\mu_{n})<\varphi\,|\,X_{n+1}=x,Z_{n+1}=z}\right),$
$B=\mathbb{P}^{\mathcal{T}}\left({f_{\tau_{x,z}}^{-1}(\lambda_{x,Y_{n+1},z})\leq\varphi\leq Q_{1-\alpha}(\mu_{n})\,|\,X_{n+1}=x,Z_{n+1}=z}\right).$
@@
$\mathbb{P}^{\mathcal{T}}\left({\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}\left({Q_{1-\alpha}(\mu_{n})}\right)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ =A+B+\mathbb{P}^{\mathcal{T}}\left({\varphi<f_{\tau_{x,z}}^{-1}\left({\lambda_{x,Y_{n+1},z}}\right)\leq Q_{1-\alpha}(\mu_{n})\,|\,X_{n+1}=x,Z_{n+1}=z}\right),$
@@
$\mathbb{P}^{\mathcal{T}}\left({\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}(\varphi)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ =A+B+\mathbb{P}^{\mathcal{T}}\left({Q_{1-\alpha}(\mu_{n})<f_{\tau_{x,z}}^{-1}\left({\lambda_{x,Y_{n+1},z}}\right)\leq\varphi\,|\,X_{n+1}=x,Z_{n+1}=z}\right).$
@@
$\mathbb{P}^{\mathcal{T}}\left({\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}\left({Q_{1-\alpha}(\mu_{n})}\right)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ -\mathbb{P}^{\mathcal{T}}\left({\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}(\varphi)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ =\mathbb{P}^{\mathcal{T}}\left({\varphi<f_{\tau_{x,z}}^{-1}\left({\lambda_{x,Y_{n+1},z}}\right)\leq Q_{1-\alpha}(\mu_{n})\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ -\mathbb{P}^{\mathcal{T}}\left({Q_{1-\alpha}(\mu_{n})<f_{\tau_{x,z}}^{-1}\left({\lambda_{x,Y_{n+1},z}}\right)\leq\varphi\,|\,X_{n+1}=x,Z_{n+1}=z}\right).$ (11)
  1. 1.
@@
    $\mathbb{P}^{\mathcal{T}}\left({\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}(\varphi)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ \geq\mathbb{P}^{\mathcal{T}}\left({\lambda_{x,\hat{Y}_{n+1},z}\leq f_{\tau_{x,z}}(\varphi)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)-\mathrm{d}_{\mathrm{TV}}(\textup{P}_{Y\mid X=x};\Pi_{Y\mid X=x}).$
@@
    $\mathbb{P}^{\mathcal{T}}\left({\lambda_{x,\hat{Y}_{n+1},z}\leq f_{\tau_{x,z}}(\varphi)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ =\mathbb{P}^{\mathcal{T}}\left({\hat{Y}_{n+1}\in\left\{{y\in\mathcal{Y}\colon\lambda_{x,y,z}\leq f_{\tau_{x,z}}(\varphi)}\right\}\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ =\mathbb{P}^{\mathcal{T}}\left({\hat{Y}_{n+1}\in\mathcal{R}_{z}(x;f_{\tau_{x,z}}(\varphi))\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ =\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau_{x,z}}(\varphi)))\geq 1-\alpha.$